The length of a rectangle is 8 cm mor than 3 times its width. the perimeter of the rectangle is 64 cm.
A. What are the dimensions of the rectangle?

Answer by checkley71(8403)

L=3W+8
PERIMETER=2L+2W
64=2(3W+8)+2W
64=6W+16+2W
64=8W+16
8W=64-16
8W=48
W=48/8
W=6 CM FOR THE WIDTH.
L=3*6+8
L=18+8
L=26 CM FOR THE LENGTH
PROOF
64=2*6+2*26
64=12+52
64=64
AREA=W*L
AREA=6*26
AREA=156 CM^2 ANSWER FOR THE AREA.